# On the isotopies of tangles in periodic 3-manifolds using finite covers

### v1

@@ Acknowledgements
The authors are grateful to Yuya Koda and Yuta Nozaki for their helpful discussions. This work is supported by the World Premier International Research Center Initiative Program, International Institute for Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2), MEXT, Japan. The first author is also supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 20K14322 and Grant-in-Aid for Scientific Research (C) Grant Number 25K07005. The second author is also supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 25K17246. The third author is also supported in part by National Science Foundation Grant No. DMR-1847172 and Research Corporation for the Advancement of Science Cottrell Scholar Award Grant No. CS-CSA-2020-162. The last author is also supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 19K14530.

### v2

@@ Acknowledgements
The authors are grateful to Kai Ishihara, Yuya Koda, and Yuta Nozaki for their helpful discussions. This work was supported by the World Premier International Research Center Initiative Program, International Institute for Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2), MEXT, Japan, and JSPS Program for Forming Japan’s Peak Research Universities (J-PEAKS) Grant Number JPJS00420230011. The first author was also supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 20K14322 and Grant-in-Aid for Scientific Research (C) Grant Number 25K07005. The second author was also supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 25K17246. The third author was also supported in part by National Science Foundation Grant No. DMR-1847172 and Research Corporation for the Advancement of Science Cottrell Scholar Award Grant No. CS-CSA-2020-162. The last author was also supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 19K14530.